OpenStudy (anonymous):

y varies inversely as three times x. When x = 3, y = 4. What is the value of x when y = 2?

hero (hero):

\[xy = k\]

hero (hero):

xy = k 3*4 = k 12 = k If y = 2 then 2x = 12 x = 12/2 x = 6